OpenStudy (anonymous):

Find the length of the length of the altitude drawn to the hypotenuse. The traingle is not drawn to scale

OpenStudy (anonymous):



OpenStudy (anonymous):

is that the only given to the problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it not stated that the triangle is 30-60-90?? i can't solved it because of lack of information

OpenStudy (anonymous):

owh.. yes.. there is an answer.. sorry

OpenStudy (ajprincess):

 Using pythagoras theorem for the triangle ABC \(x^2+y^2={31}^2\) Using pythagoras theorem for the triangle ABD \(x^2=5^2+z^2\)-(1) Using pythagoras theorem for the triangle BDC \(y^2=z^2+{26}^2\)-(2) Nw add (1) and (2). then u will get (1)+(2)=>\(x^2+y^2=5^2+z^2+z^2+{26}^2\) \({31}^2=5^2+z^2+z^2+{26}^2\) \(961=25+2z^2+676\) Find z. Can u do?